Circular No. 7716

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE R. Chornock, W. D. Li, and A. V. Filippenko, University of California

at Berkeley, report that inspection of CCD spectra (range 330–1000 nm) obtained on Sept. 12 UT with the Shane 3-m reflector at Lick Observatory reveals that SN 2001ef (IAUC 7710) is a supernova of type Ic, not type Ia (cf. IAUC 7714); the spectrum resembles that of SN 1994I (Filippenko 1997, Ann. Rev. Astron. Astrophys. 35, 309), about a week past maximum brightness. Spectra obtained as above show that SN 2001dn (IAUC 7681) is a supernova of type Ia, about a month past maximum brightness. SN 2001dt (IAUC 7689), SN 2001dw (IAUC 7691), and SN 2001ec (IAUC 7702) are type-Ia supernovae, several weeks past maximum brightness. Superposed narrow emission lines of the host galaxy of SN 2001ec are at redshift cz ≈ 14 000 km/s. SN 2001eg (IAUC 7712) is type-Ia supernova, several months past maximum brightness. SN 2001dy (IAUC 7693) is a type-II supernova, a few weeks past maximum brightness. SN 2001dz (IAUC 7696) and SN 2001ea (IAUC 7696) are supernovae of type II, probably not long after maximum light, although the Hα emission component substantially dominates over the absorption. SN 2001ee (IAUC 7705) is a type-II supernova several months past maximum light. Chornock and Filippenko add that a CCD spectrum (range 330–1000 nm) of SN 2001dx (IAUC 7692) obtained on Sept. 11 reveals that the object is of type II, with the usual P-Cyg profiles of hydrogen.

T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrum of SN 2001ec, obtained by P. Berlind on Sept. 12.44 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-I supernova, but it is unclear if it is of type Ic near maximum light or type Ia several weeks past maximum light. Narrow Hα emission from the nucleus indicates a recession velocity of 13 600 km/s for the host galaxy.

SUPERNOVA 2001du IN NGC 1365 A. J. Beasley and J. M. Carpenter, Owens Valley Radio Observatory,

California Institute of Technology, write: “Radio observations of SN 2001du (IAUC 7690, 7692, 7704, 7705) using the Owens Valley Millimeter Array at 90 GHz on Aug. 27 (five antennas) and Sept. 4 (four antennas) place 3σ upper limits of 11 and 14 mJy (respectively) on any high-frequency radio emission associated with the supernova.”

2001 September 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green